### v2

# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

Abstract
In the domain of Multimodal Large Language Models (MLLMs), achieving human-centric video understanding remains a formidable challenge. Existing benchmarks primarily emphasize object and action recognition, often neglecting the intricate nuances of human emotions, behaviors, and speech-visual alignment within video content. We present HumanVBench, an innovative benchmark meticulously crafted to bridge these gaps in the evaluation of video MLLMs. HumanVBench comprises 16 carefully designed tasks that explore two primary dimensions: inner emotion and outer manifestations, spanning static and dynamic, basic and complex, as well as single-modal and cross-modal aspects. With two advanced automated pipelines for video annotation and distractor-included QA generation, HumanVBench utilizes diverse state-of-the-art (SOTA) techniques to streamline benchmark data synthesis and quality assessment, minimizing human annotation dependency tailored to human-centric multimodal attributes. A comprehensive evaluation across 22 SOTA video MLLMs reveals notable limitations in current performance, especially in cross-modal and emotion perception, underscoring the necessity for further refinement toward achieving more human-like understanding. HumanVBench is open-sourced to facilitate future advancements and real-world applications in video MLLMs.
Figure 1: Overview of HumanVBench, which encompasses 16 fine-grained tasks for extensive human-centric evaluations (middle blue box). Each task is denoted by its acronym and the number of included QA instances. The right orange box illustrates some examples of these QAs. HumanVBench is constructed using the novel automated Video Annotation Pipeline (upper left, purple box), followed by the Distractor-Included QA Synthesis Pipeline (lower left, green box). These pipelines are reusable and backed by more than twenty data processing operators with advanced algorithm implementation and cutting-edge auxiliary models.
1 Introduction
In recent years, Multimodal Large Language Models (MLLMs) have emerged as pivotal technological advancements, significantly expanding traditional language model capabilities to include processing and comprehending diverse data forms like text, images, and videos [44, 8, 3]. Among these, video-oriented MLLMs [55, 11, 34] have garnered substantial research interest due to their potential in interpreting video content in a manner closely aligned with human perception. While image-based MLLMs [36, 38, 62, 50, 49, 24] primarily focus on static content, video-MLLMs offer enhanced capacities for understanding the complex temporal dynamics intrinsic to video data.
Human-centric scenes in videos naturally attract attention due to the emphasis on individuals’ emotions, actions, and verbal interactions, necessitating effective comprehension by video understanding models. Despite advances in this field, existing benchmarks often fall short in rigorously assessing the nuanced understanding of human emotions and behaviors. Current evaluations predominantly focus on general content comprehension, object recognition, and action detection, often neglecting subtle intricacies such as emotional insight and behavioral analysis. Furthermore, synchronizing speech with visual elements remains a substantial challenge; unlike humans, who effortlessly discern mismatches between audio and visual cues, computational models often struggle with tasks like identifying speakers and aligning speech with corresponding lip movements.
To bridge these gaps, we introduce HumanVBench, a pioneering benchmark specifically tailored for video MLLMs focusing on human-centric analysis. HumanVBench includes 16 meticulously designed QA tasks, categorized into two core aspects: inner emotion and outer manifestation, as illustrated in Figure 1. The inner emotion dimension evaluates the model’s capacity to perceive emotional cues from videos, while the outer manifestation encompasses person recognition, human behavior analysis, and cross-modal speech-visual alignment. Each category is further subdivided into many fine-grained tasks, offering a comprehensive evaluation across static and dynamic, basic and complex, single-modal and cross-modal aspects. For instance, it enables us to examine MLLMs’ understanding of human actions, emotions, and the associated causal relationships, as well as consider the alignment degree between appearance, mouth movement, audio, and speech content in videos.
The construction of HumanVBench is facilitated by the advanced open-source data processing system, Data-Juicer [6], and involves two novel data synthesis pipelines: the Human-Centric Video Annotation Pipeline and the Distractor-Included QA Synthesis Pipeline, leveraging over twenty state-of-the-art (SOTA) data processing operators. Unlike conventional benchmarks that rely heavily on human annotators [26, 39, 42, 51, 17, 18], our approach automates multi-modal, fine-grained annotation processes, with the help of cooperation by diverse task-specific annotation algorithms and models. The Distractor-Included QA Synthesis Pipeline uses iterative enhancement with video-MLLMs to refine question accuracy and generate distractors, while ensuring quality through a final human review. Thanks to the fruitful annotations and dedicated orchestrations of data processing operators, the construction of HumanVBench is largely automated, substantially reducing the need for manual intervention. Additionally, our method is applicable to “in-the-wild” video data, enabling the creation of video benchmarks that are not confined to controlled or domain-specific environments.
Through HumanVBench, we comprehensively evaluate 22 SOTA video MLLMs, including open-source models like VideoLLaMA3 [57] and commercial ones like GPT-4o [41]. Our evaluations reveal several interesting insights and significant gaps between current model capabilities and human-like understanding, particularly in tasks such as cross-modal alignment and emotion perception. While proprietary models demonstrate closer human-like accuracy, open-source models frequently misclassify emotions due to temporal noise, underscoring the need for further architectural improvements and refined datasets.
In summary, our contributions are as follows:
  • We introduce HumanVBench, a novel video benchmark for MLLMs that emphasizes fine-grained human comprehension in videos, focusing on emotion perception, person identification, behavioral analysis, and speech-visual alignment.
  • We propose two advanced pipelines for automatic video annotation and the production of high-quality, multiple-choice questions relevant to video-based descriptive queries. These pipelines reduce manual labor and are equipped with diverse reusable operators for multimodal, detailed individual labeling in videos, making them adaptable across various contexts.
  • Our comprehensive evaluation of numerous SOTA video MLLMs offers key insights, facilitating in-depth discussions regarding their performance, strengths, and areas for enhancement.
  • We release our benchmark, including data, evaluation, and synthesis codes at https://github.com/modelscope/data-juicer/tree/HumanVBench , to foster further evolution of future human-centric video analysis systems.
2 Related Works
Multimodal Large Language Models. The remarkable progress in Large Language Models (LLMs) has sparked extensive research into merging language comprehension with visual and auditory information, thereby expediting the advancement of multimodal models. Within this domain, image MLLMs amalgamate visual and linguistic data to enhance image interpretation and cross-modal reasoning [13, 32, 23, 37]. Video MLLMs [29, 39, 34, 48, 58, 16] extend these capabilities by incorporating temporal sequences for dynamic video analysis. Furthermore, generalist MLLMs [20, 19, 61] can process a plethora of inputs, including images, video, and audio, thus improving adaptive task performance across various modalities. Despite these advancements, rigorous evaluation of video MLLMs on human-centric video understanding tasks remains an unaddressed challenge, which this work aims to address.
Video Benchmarks for MLLMs. There exists a substantial array of benchmarks for assessing the performance of MLLMs, with a particular emphasis on video-based evaluations. Presently, general-purpose video benchmarks can be broadly categorized based on their evaluative methodologies: multiple-choice queries with definitive answers [40, 42, 30, 18, 45], Open-Ended Video Question Answering necessitating supplementary LLMs for evaluation [52, 39, 17, 45], video-captioning benchmarks [52, 12, 5], and video-generation assessments such as VBench [22]. Additionally, there are human-centric benchmarks in MLLMs. ActivityNet-QA [52], rooted in the widely-used ActivityNet dataset [56], emphasizes object and human activity recognition. HumanBench [46] amalgamates multiple public datasets to create benchmarks for tasks like pose estimation, pedestrian attribute recognition, and crowd assessment. HERM [31] introduces a human-centric image benchmark focusing on people, poses, actions, and interactions, using GPT-4-generated question-answer pairs.
Our Position. Compared to existing benchmarks, HumanVBench distinguishes itself through both innovative task design and unique construction methodologies. Firstly, it introduces novel human-related video tasks. Unlike HERM, HumanVBench integrates complex scenarios encompassing temporal dynamics for enhanced human video interpretation; compared to ActivityNet-QA and HumanBench, HumanVBench delves into nuanced emotional comprehension and evaluates cross-modal alignment between visual data and speech modalities. Moreover, HumanVBench boasts a pioneering construction approach, deriving from raw, uncurated video content and employing over 20 advanced processing operators for meticulous video character annotation and automatized multiple-choice question formulation. Human intervention is minimized to verify the quality of well-structured QAs.
3 The Proposed HumanVBench
3.1 Task Design and Definition
Human observers naturally concentrate on individuals in videos, examining their appearance, emotions, and behaviors. This intrinsic focus underpins our design of 16 fine-grained, human-centric tasks, aimed at evaluating MLLMs’ ability to mimic human-like perception and understanding in video analysis (Figure 1). Each task is detailed with definitions and examples in the Appendix Section 10. The tasks are grouped into two categories based on content observability: Inner Emotion and Outer Manifestation.
3.1.1 Inner Emotion
In real-world videos, inner emotions are less observable, as cameras often capture wider scenes instead of facial close-ups. Thus, emotion perception becomes an abstract skill requiring careful attention to facial expressions, body language, and verbal cues. With this goal, tasks in this category evaluate video-MLLMs’ ability to detect and interpret emotional nuances, mirroring human perceptual skills.
Specifically, five specific tasks comprise the Emotion Perception category:
  • Emotion Recognition (ER) requires identifying the most fitting emotional description for an individual in a video.
  • Emotion Temporal Analysis (ETA) focuses on tracking emotional changes over time.
  • Attitude Recognition (AT) assesses the inferred attitudes of individuals in relation to specific events or entities, classifying them as positive, negative, or neutral.
  • Emotion Intensity Comparison (EIC) evaluates the model’s ability to differentiate and quantify the emotional intensity of various individuals.
These tasks collectively enable the evaluation of MLLMs’ potential in capturing emotional cues and facial details from videos, thereby gauging their capabilities in inner emotional understanding.
Figure 2: The Human-Centric Video Annotation Pipeline involves obtaining videos featuring people and annotating both visual and auditory information as well as overall event atmospheres.
Figure 3: The Distractor-Included QA Synthesis Pipeline facilitates four steps: selecting “question videos”, generating preliminary answers, optimizing answers with generated distractors, and manually verifying multiple-choice questions.
3.1.2 Outer Manifestation
Unlike inner emotions, outer manifestations deal with more tangible aspects such as identifying individual(s), causality reasoning, and synchronization of video elements like speech or singing. Guided by these aspects of human observation, we formulated three task categories:
The first basic dimension is for Person Recognition, evaluating the model’s capability to identify a particular person within a complex scene, akin to “person-finding”:
  • Text-to-Human (T2H) tests the model’s ability to identify a person based on textual descriptions.
  • Human-to-Text (H2T) assesses the accuracy of text explanations attributed to a target person, distinguishing them from others.
  • Human Counting (HC) evaluates the model’s ability to detect, track, and count distinct individuals in a video.
  • Appearance Time Detection (ATD) requires identifying a specified individual’s presence time in the video.
Besides, we incorporate the Human Behavior Analysis category to scrutinize the model’s ability to understand and analyze individual behaviors, with four tasks:
  • Behavior Temporal Analysis (BTA) explores behavior tracking capabilities over time.
  • Behavior Causality Analysis (BCA) examines causal inferences within behavior sequences.
  • Action at Specified Time (AST) requires identifying precise actions at given times.
  • Time of Specific Action (TSA) tests the detection of exact moments specifying actions occur.
Furthermore, a critical cross-modal category is Speech-Visual Alignment, designed to test capabilities involving the following tasks:
  • Audio-Visual Speaker Matching (AVSM) correlates audio features to identify individuals and analyzes matches by appearance (gender and age).
  • Active Speaker Detection (ASD) identifies the individual currently speaking by integrating visual and audio inputs.
  • Audio-Visual Alignment Detection (AVAD) detects synchronization points, examining the coherence between lip movements and audio.
  • Speech Content Matching (SCM) requires analyzing spoken content against text, and evaluating transcription or lip-reading capabilities.
These aforementioned tasks together enable a comprehensive assessment of MLLMs’ diverse perceptual abilities across emotion, identity, behavior and speech-visual alignment, pushing toward human-like video comprehension.
3.2 Human-Centric Video Annotation Pipeline
Creating task-specific questions for the aforementioned tasks hinges on extensive human-centric annotations within videos. Our video annotation pipeline, illustrated in Figure 2, emphasizes multi-modal, granular annotation pathways, supported by the Data-Juicer [6, 9] framework. While benefiting from existing operators, we’ve also developed innovative operators to enhance the open-source community. Detailed annotations and examples are in the Appendix Section 11.
3.2.1 Collecting Videos Containing People
We sourced copyright-free videos from Pexels [43], splitting each video by scene transitions using video_split_by_scene_mapper for accurate human tracking. After a series of attribute-based filtering such as duration and optical flow, a video_face_ratio_filter gauges the visibility of faces, retaining videos where a face appears in most frames for further subsequent annotation.
3.2.2 Video Mappers
We begin by implementing a specially designed operator named video_human_tracks_extraction_mapper to track each person across consecutive frames. The operator constructs tracks by iteratively detecting face and body bounding boxes, using overlap thresholds between consecutive frames to ensure accurate and consistent person localization. Additionally, this process enables an approximate count of the number of people appearing within the continuous shot. The generated tracks serve as a foundation for subsequent tasks, including person highlighting and detailed descriptions of their appearances or actions.
Next, visual attributes and descriptions for each individual are derived from track information. Specifically, the human_demographics_mapper extracts facial images from face bounding box tracks and employs existing facial attribute detection models to infer demographic labels (e.g., age, gender, race). Additionally, bounding box data at various positions (e.g., full-body or face-only) enables flexible cropping to generate individual-focused videos. This allows the description model to concentrate on appearance and posture (via the video_human_description_mapper) or facial expressions and transformations (via the video_facial_description_mapper). These processes effectively prevent the description model from misidentifying bounding box contexts or being disrupted by other individuals in the frame, while ensuring facial details are preserved during resizing.
Further, audio operations enrich video content by analyzing people-related information. The annotation process initiates with the audio_tagging_mapper classifying sound types. For speech-detected videos, specialized operators are activated: the active_speaker_detection_mapper identifies active speakers by fusing audio and visual cues, the asr_mapper transcribes speech content, the speech_emotion_recognition_mapper extracts emotional cues, and the voice_demographics_mapper profiles voice characteristics (e.g., gender, age).
Finally, annotations are extended to capture event atmosphere and broader video narratives through the video_description_mapper. These detailed annotations facilitate versatile, automated question synthesis. For example, in the Audio-Visual Speaker Matching task construction, videos are filtered by verifying that exactly one individual’s Label-3 (visual demographics) matches the video’s Label-10 (audio demographics), with all other individuals exhibiting mismatches. The matching individual is established as the ground truth, with others marked as interference, while Label-1 tracking data assigns distinct identifiers to each person. Complete task construction details are provided in the Appendix Section 12.
3.3 Distractor-Included QA Generation Pipeline
For tasks with exact answers (e.g., restricted categories, numbers, or letters), questions, correct answers, and distractors are constructed using tailored templates and annotations from Section 3.2. For open-ended questions, we developed a pipeline to generate questions, answers, and distractors, applied to six tasks in HumanVBench: Human Emotion Recognition, Emotion Temporal Analysis, Emotion Intensity Comparison, Human-to-Text, Behavior Temporal Analysis, and Behavior Causality Analysis. The Distractor-Included QA Synthesis Pipeline is illustrated in Figure 3.
3.3.1 Selecting Videos for Generating Questions
For a specific task, annotated videos are filtered based on task-specific criteria, such as video length and the number of people, to select suitable candidates for question generation. For instance, the Behavior Temporal Analysis task requires longer videos, while Human-to-Text necessitates at least two people. This process yields a subset of videos tailored for question generation.
Human Emotion Perception Person Recognition Human Behavior Analysis Speech-Visual Alignment
Models Frames ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25 25.0 25.0 25.0 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.0 31.2
Chat-UniVi 1 f/s 29.2 19.9 10.9 16.3 18.8 29.1 37.2 17.8 19.8 26.0 22.0 19.4 14.0 6.7 15.5 42.8 20.6 26.3 18.6 27.1
CogVLM2-Video 1 f/s 25.0 33.1 36.0 11.1 25.8 42.5 63.8 31.1 40.0 44.4 42.4 47.8 34.0 14.1 34.6 59.2 38.7 30.7 17.9 36.6
VideoLLaMA3 1 f/s 39.4 34.5 48.0 43.1 41.3 90.3 78.1 44.9 71.0 71.1 72.4 53.4 56.4 69.0 62.8 64.6 63.1 34.4 17.9 45.0
VILA 6 f 29.1 34.4 28.7 14.7 26.7 50.4 54.2 40.4 20.2 41.3 53.0 54.5 33.0 51.4 48.0 47.4 23.9 34.3 18.6 31.1
Video-LLaVA 8 f 27.1 25.7 26.2 24.7 25.4 27.9 46.8 28.3 31.7 33.7 34.8 37.3 27.0 40.7 35.0 50.0 28.4 34.3 21.2 33.5
LLaVAOneVision 8 f 36.5 33.1 62.8 15.7 36.0 67.0 68.1 49.1 10.0 48.6 63.6 55.2 37.0 47.5 50.8 52.6 51.6 35.0 26.9 42.8
InternVL2 8 f 35.4 29.3 40.2 25.9 32.7 70.4 59.6 37.2 20.8 47.0 58.5 52.2 38.0 33.8 45.6 51.7 55.0 33.6 25.5 41.5
InternVL2.5 8 f 44.8 31.6 54.3 51.5 45.6 81.0 78.7 40.6 35.8 59.0 72.9 55.2 41.0 62.1 57.8 65.1 61.3 32.1 15.4 43.5
Qwen-VL2 8 f 41.7 40.4 42.7 32.8 39.1 79.3 69.2 43.4 20.8 53.2 61.9 55.2 32.0 51.4 50.1 50.7 56.1 31.4 23.7 40.5
Qwen-VL2.5 8 f 43.7 31.6 35.4 46.5 39.3 88.8 77.6 50.9 30.8 62.0 72.9 47.8 34.0 50.8 51.4 71.0 61.3 33.6 18.8 46.2
PLLaVA 16 f 25.0 24.3 36.6 23.2 26.5 33.5 54.3 29.9 23.3 35.3 30.5 32.8 34.0 14.7 28.0 50.7 37.4 29.9 21.8 35.0
ShareGPT4Video 16 f 36.5 30.9 40.9 10.1 29.6 33.0 40.4 24.5 43.3 35.3 45.8 40.3 39.0 16.9 35.5 44.1 31.0 34.3 25.0 33.6
Otter-V 16 f 15.6 25.7 26.2 21.2 21.9 30.2 22.3 20.8 22.5 24.0 22.9 17.9 20.0 23.1 21.0 38.2 26.5 33.6 27.6 31.5
VideoChat2-IT 16 f 25.0 30.9 39.6 25.3 30.2 20.1 47.9 11.2 26.7 26.5 42.4 47.8 34.0 24.9 37.3 43.4 27.7 31.4 23.0 31.4
LLaVA-Video 64 f 40.6 32.4 59.8 37.4 41.8 74.3 72.3 44.3 26.7 54.4 68.6 52.2 47.0 58.8 56.7 52.0 58.7 32.8 39.3 45.7
Video-LLaMA 8 f 25.0 25.9 11.5 19.7 18.9 28.4 33.2 23.9 20.0 26.4 39.5 32.5 23.2 15.4 27.7 40.1 26.6 33.1 26.2 31.5
VideoLLaMA2.1-AV 8 f 32.3 28.7 47.0 21.7 31.9 34.6 47.9 41.5 16.7 35.2 47.5 43.3 27.0 33.3 37.8 44.0 31.6 32.1 23.7 32.9
ImageBind-LLM 15 f 15.6 21.0 25.0 26.4 21.0 23.5 24.5 23.8 19.5 22.8 19.5 19.4 24.0 23.3 21.6 45.0 25.1 28.9 22.9 30.5
ChatBridge 4 f 27.1 15.6 30.2 14.4 21.8 31.4 41.5 23.7 8.7 26.3 31.4 23.9 41.0 16.9 28.3 43.7 25.3 30.4 24.4 30.9
OneLLM 15 f 26.0 29.4 34.1 21.1 26.1 29.0 38.3 20.6 27.8 28.9 32.2 38.8 23.0 22.4 29.1 43.4 26.4 29.5 23.2 30.6
GPT-4o 54.1 38.2 27.4 64.1 40.1 46.9 83.0 47.2 32.5 50.7 81.4 67.2 47.1 61.6 64.3 - - - - -
Gemini-1.5-Pro 57.3 54.4 53.0 67.2 55.9 87.1 77.7 52.8 71.7 72.3 78.0 65.7 54.0 75.1 68.2 90.1 76.8 66.4 84.6 79.5
Human 91.6 86.0 87.8 80.8 86.6 98.9 85.8 92.5 78.3 88.9 93.0 86.5 88.6 88.1 89.1 96.0 96.1 87.0 88.0 91.8
Table 1: The performance of SOTA video MLLMs on HumanVBench, grouped into 15 visual-only MLLMs, 5 audio-visual MLLMs and 2 proprietary ones. “Random” denotes random guessing, and “Human” indicates human-level performance. Task acronyms are defined in Figure 1. For each task, the best overall result is bolded, and the best open-source result is underlined. “-” means the model recognizes its lack of required capabilities for the task and thus refuses to answer. The results for all open-source models are averaged over five runs with different random seeds. Additional evaluation details are in the Appendix Section 9.
3.3.2 Question and Preliminary Answer Synthesis
This stage synthesizes questions and initial answers by processing the video with bounding box track information of the target person. For tasks focusing on facial expressions (Human Emotion Recognition and Emotion Temporal Analysis), we reconstruct the video using cropped face regions; for others, we add a red bounding box to create “marked videos”. Then, we design prompts for Video-MLLM to generate task-specific captions centered on the target individual. We find that video-MLLMs don’t always focus correctly on the person highlighted by the bounding box. To address this, we enhance attention by adding a visual cue of the target person’s appearance from Label-4 (Figure 2).
After acquiring task-specific captions, we analyze the caption attribute distributions (e.g., sentiment distribution for emotion recognition) and filter materials for question generation based on the desired distribution. Using these captions, we leverage SOTA LLMs to formulate questions and preliminary answers according to task definitions.
3.3.3 Answer Optimization and Distractors Generation
To obtain the most suitable answer, we iteratively optimize the answer using three Video-MLLMs while generating three distractors. Each Video-MLLM first generates a candidate answer from the video and question, then selects a better answer between this candidate and the current optimal answer (initialized as the preceding preliminary answer), updating the optimal answer. The discarded answers serve as raw material for distractor generation. This process repeats three times, producing one optimal answer and three discarded answers. An LLM then modifies the discarded answers with task-specific disturbances, ensuring distractors are distinct from the correct answer.
By using MLLM-discarded answers to synthesize options, the distractors are more challenging compared to those directly generated from the question and answer, thus more rigorously testing the model’s discriminative ability.
3.3.4 Manual Verification and Correction
Given the limitations of the annotation and question-answer generation models, errors in the multiple-choice questions may occur. To address this, we conduct manual validation as a final step. Specifically, we shuffle the options of each generated question and ask verifiers to select the best answer. If their choice matches the pipeline’s automatic selection, we confirm the question aligns with human cognition. If not, we re-examine and determine whether to revise the correct answer. If no suitable option is found, verifiers provide the correct answer, which becomes the definitive answer for the question. While this process involves some manual effort, it is significantly less labor-intensive and costly than generating questions and answers from scratch.
The Distractor-Included QA Synthesis Pipeline automates the tedious process of creating questions and options, shifting the human role from question creators to quality inspectors. Notably, we’ve decomposed the generation process into simpler sub-tasks, each manageable by open-source (M)LLMs. While these models are less accurate than their commercial counterparts, their inherent noise enhances the cognitive complexity of distractors. After human review and correction, this approach can still ensure high question quality while reducing labor costs. The pipeline is highly transferable, enabling its adaptation for designing a broader range of descriptive questions and choices.
3.4 Post-Processing
To address the prevalent issue of answer leakage [10] in multimodal evaluation datasets, we adopt the approach proposed in [10]. Specifically, we test all evaluated models without visual input and remove frequently correct QAs (approximately 6%) to ensure random accuracy. This strategy preserves the visual relevance of the questions while effectively mitigating the risk of answer leakage. Finally, HumanVBench includes 2116 problem instances. Instance counts for each task are detailed in Figure 1, and additional statistics are available in Appendix Section 7.
Figure 4: Two examples of 8-frame speaker videos sampled at equal intervals in the emotion recognition task, along with the responses from different MLLMs.
Models Chat-UniVi LLaVA-OneVision Intern VL2 Share-GPT4-Video Video Chat2-IT Qwen-VL2 LLaVA-Video Video-LLaMA Video-LLaMA-2.1 Chat Bridge GPT-4o Gemini-1.5-Pro Avg. Dec.
Emo. Acc. 29.2 36.5 35.4 36.5 25.0 41.7 40.6 25.0 32.3 27.1 54.1 57.3 36.7
Speaker Emo. Acc. 24.3$\downarrow$-4.9 34.3$\downarrow$-2.2 27.1$\downarrow$-8.3 32.8$\downarrow$-3.7 21.4$\downarrow$-3.6 37.1$\downarrow$-4.6 37.1$\downarrow$-3.5 22.9$\downarrow$-2.1 31.4$\downarrow$-0.9 21.4$\downarrow$-5.7 51.6$\downarrow$-2.5 54.0$\downarrow$-3.3 32.9$\downarrow$-3.8
Table 2: Accuracy of Video-MLLMs on Emotion Recognition (ER) for the full dataset and the subset where target individuals are in the speaking state. Current video-MLLMs struggle with emotion recognition for speaking individuals in HumanVBench.
4 Evaluation and Insights
4.1 Experimental Settings
We meticulously select 20 SOTA open-source video MLLMs. These included both visual-only models such as Chat-UniVi [25], CogVLM2-video [21], LLaVA-One-Vision [28], PLLaVA-7b [53], ShareGPT4Video [11], Otter-V [27], VILA [35], VideoChat [29], InternVL series [14, 15], Qwen-VL series [48, 4] and VideoLLaMA3 [57] and audio-visual models capable of analyzing both visual and audio inputs, such as the Video-LLaMA series [58, 16] and generalist MLLMs like ImageBind-LLM [19], ChatBridge [61], and OneLLM [20]. We also evaluated commercial models GPT-4o [41] and Gemini-1.5-Pro [47]. All QAs were framed as multiple-choice questions (N choose 1, with N varying across different test samples), reporting both accuracy and the performance of random guesses and graduate-level humans for reference. More implementation details can be found in the Appendix Section 9.
4.2 Main Results
Table 1 summarizes the performance evaluation results for our benchmark, revealing several key insights from analyzing vision-only tasks (emotion perception, person recognition, and behavior analysis tasks) and cross-modal tasks (speech-visual alignment tasks).
4.2.1 Performance in Vision-Only Tasks
Open-Source Video-MLLMs: Generally, these models still show a noticeable gap compared to human-level performance, particularly in Emotion Perception tasks, highlighting the need for further development to bridge this gap. However, some models have demonstrated superior performance in certain tasks, rivaling or even surpassing commercial models, such as LLaVA-OneVision in AR and VideoLLaMA3 in T2H, ATD, and AST. Among open-source models, VideoLLaMA3 leads in video human understanding, and achieves a modest 58.4% mean accuracy across 12 vision-only tasks. Although it still lags behind human performance (88.2%), it has surpassed the commercial model GPT-4o (51.7%) and is close to Gemini-1.5-pro (65.5%), demonstrating the immense potential of open-source models.
Proprietary MLLMs: Overall, these models outperform their open-source counterparts in most tasks, achieving near-human proficiency in several tasks. Notably, Gemini delivers superior performance compared to other models, while GPT-4o struggles in tasks like AR, T2H, and ATD. As a result, GPT-4o’s average performance in Emotion Perception and Person Recognition lags behind several leading open-source models.
4.2.2 Performance in Speech-Visual Alignment Tasks
Lip-Reading Ability of Visual-Only MLLMs For AVSM and ASD tasks, many videos include dubbing or multi-speaker conversations with single-speaker audio, making lip movements alone insufficient for accurate responses. Therefore, when answering these two tasks, visual-only models essentially degrade to speech action recognition but can still get some questions correct. As shown in Table 1, these MLLMs show some lip-reading ability, with QwenVL2 and InternVL2 outperforming other models. However, for the SCM task, almost all models perform at a random level, indicating that current models lack precise lip-reading (lip translation) ability. Future models could focus on enhancing lip-reading capabilities.
Error Type Error Percentage
Misidentifying emotion as Surprise 41%
Overinterpretation of neutral emotions 29%
Emotion Polarity Mistakes 29%
Table 3: Error Analysis of Emotion Recognition in Gemini
Video-MLLMs ATD AST TSA
Chat-UniVi 17.5 ($\uparrow$2.3) 40 ($\uparrow$26) 14.1 ($\uparrow$7.4)
CogVLM2 40.8 ($\uparrow$0.8) 35 ($\downarrow$1) 10.7 ($\downarrow$3.4)
VILA 38.3 ($\uparrow$18.1) 33 (–) 48 ($\downarrow$3.4)
Video-LLaVA 36.7 ($\uparrow$5) 25 ($\downarrow$2) 45.8 ($\uparrow$5.1)
LLaVAOneVision 10 (–) 41 ($\uparrow$4) 47.5 (–)
InternVL2 31.7 ($\uparrow$10.9) 45 ($\uparrow$7) 52.5 ($\uparrow$18.7)
Video-LLaMA 24.2 ($\uparrow$4.2) 25 ($\uparrow$7) 9 (–)
Video-LLaMA2.1 24.2 ($\uparrow$7.5) 25 ($\downarrow$2) 29.4 ($\downarrow$3.9)
ChatBridge 17.5 ($\uparrow$8.8) 40 ($\downarrow$1) 14.1 ($\downarrow$2.8)
Average 26.8 ($\uparrow$5.9) 34.3 ($\uparrow$4.4) 30.1 ($\uparrow$2.0)
Table 4: Timestamp integration effect on Video-MLLMs in Appearance Time Detection (ATD), Action at Specific Time (AST) and Time of Specific Action (TSA).
Audio-Video MLLMs: Despite handling both audio and visual inputs, these open-source MLLMs perform near-randomly across Speech-Visual Alignment tasks, particularly in AVSM and ASD, where they lag behind visual-only MLLMs. On one hand, this may stem from deficiencies in the models’ ability to visually interpret lip movements (see Appendix Section 8 for extended experiments); on the other, contemporary open-source models largely neglect intricate speech-to-lip movement alignment. Few have explicit architectural encoding linking audio and video modalities to facilitate such synergy, compounded by a scarcity of comprehensive datasets proposing coherent visual-audio lexical mapping. Consequently, these models struggle acutely in correlating speech to lip movements, though dedicated ASR models effectively address Speech Content Matching — a capability absent in tested open-source video-MLLMs.
4.3 Discussion on Speaker Emotion Recognition
A detailed analysis reveals that models frequently misattribute speaker emotions as “surprise” or “shock.” For example, Gemini-pro-1.5, the top performer in Emotion Recognition, exhibits this tendency, with surprise misidentification accounting for 40% of its errors (Table 2). This issue primarily stems from frame sampling processes, where frames depicting “mouth-opening” motions are misclassified. Figure 4 illustrates this phenomenon using the commonly adopted eight-frame fixed sampling approach [34, 60, 28, 58], which introduces temporal noise and leads to erroneous emotional judgments. The speaker-centric evaluations presented in Table 2 further confirm a consistent decline in accuracy across all models, underscoring the inherent challenges of this task for video MLLMs.
4.4 Impact of Timestamp Inclusion on Time-Specific Tasks
Table 1 shows that many open-source video-MLLMs struggle with time-specific tasks—like Appearance Time Detection, Action at Specific Time, and Time of Specific Action—due to challenges in establishing “video scene-event time” correlations. Analysis shows that only a few models, including top performers like VideoLLaMA3 and LLaVA-Video, explicitly incorporate textual timestamps as part of their built-in prompt processing from sampled video frames. To explore the impact of timestamp integration, we added timestamps (e.g., “You have read N frames, corresponding to seconds in the video as: […]”) to models without native support. As shown in Table 4, this intervention generally improved temporal reasoning accuracy, though its effectiveness varied across models, likely due to limited exposure to timestamp-related data, reflecting the models’ inherent capabilities and potential.
Figure 5: Effectiveness in generating multiple-choice questions for six descriptive question types through the Distractor-Included QA Synthesis Pipeline.
4.5 Availability of Question Generation
The generation pipeline’s efficacy for multiple-choice questions across six descriptive categories is visualized in Figure 5. On average, 30% aligned directly with human responses, 20% included correct options conflicting with automated answers, and 50% required manual verification. While video-MLLMs and LLMs currently exhibit limitations in question generation, prospective advancements promise to enhance question accuracy, minimizing human intervention. Moreover, utilizing flawed video-MLLMs to craft misleading options augments evaluative complexity, thus refining model discriminative testing rigor.
5 Conclusion
We present HumanVBench to address the pressing need for improved assessment of human-centric video understanding in MLLMs. By incorporating extensive evaluation dimensions through 16 fine-grained tasks, HumanVBench provides a systemic view into both successes and critical shortcomings of video MLLMs, particularly in emotion perception and speech-visual alignment. Experimental findings across over twenty leading video MLLMs illustrate that while proprietary models approach human accuracy in some tasks, substantial advancements are required, particularly in cross-modal domains where alignment between speech and visual elements proves challenging.
Our work opens several avenues for future research. These include expanding the benchmark to encompass broader scenarios, such as video generation tasks; enhancing model architectures to improve temporal and multimodal fusion; and refining datasets to more effectively capture the nuances of human emotions and intentions. By open-sourcing the benchmark and underlying methodologies, we hope HumanVBench can foster collaborative efforts aimed at advancing the frontiers of human-like video understanding capabilities in MLLMs.
Supplementary Material
6 Overview
In the appendix, we first We provide more benchmark statistics in Section 7, then the modality ablation experiments in VideoLLaMA2 in Section 8, followed by additional evaluation details in Section 9, followed by the detailed definition and examples for the 16 tasks of HumanVBench in Section 10. Then, we present implementation specifics of each operator for the proposed Human-Centric Annotation Pipeline, illustrated by an example of the annotation process in Section 11. Finally, we introduce the construction details of all tasks and the work of the human annotators in Section 12.
7 More Statistics of HumanVBench
HumanVBench focuses on short video understanding, specifically videos with a duration of 10 seconds or less. It includes a total of 2116 question instances, with the specific number for each task indicated in 1. The total video duration amounts to 4.7 hours and demonstrates a variety of people, scenes, and video shooting styles, as shown in 6.
Figure 6: The distribution of the number of people, scenes, and video shooting styles in HumanVBench
8 Modality Ablation in VideoLLaMA2
Models Input Modal Human Emotion Perception Person Recognition Human Behavior Analysis Speech-Visual Alignment
ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25 25.0 25.0 25.0 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.0 31.2
Video-LLaMA2.1-AV A, V 32.3 28.7 47.0 21.7 32.4 34.6 47.9 41.5 16.7 35.2 47.5 43.3 27.0 33.3 37.8 44.0 31.6 32.1 23.7 32.9
Video-LLaMA2.1-AV V 27.1 31.6 45.1 17.2 30.3 36.9 46.8 36.8 15.0 33.9 43.2 50.7 31.0 29.9 38.7 43.4 29 32.8 18.8 31.0
VideoLLaMA2-7B-16F V 32.3 30.9 34.8 23.7 30.4 37.4 51.1 36.8 15.8 35.3 53.4 50.7 32.0 49.2 46.3 47.4 38.1 32.1 15.4 33.3
Table 5: The performance of VideoLLaMA2’s vision-only version (VideoLLaMA2-7B-16F), and the audio-visual version (VideoLLaMA2.1-AV), on HumanVBench, based on different modal inputs (A for Audio, V for Visual).
Despite audio-visual MLLMs processing audio data, they perform at random-guess levels on AVSM and ASD tasks, underperforming relative to many vision-only models that rely solely on lip movement analysis. This raises the question: does the poor performance stem from limitations in visual analysis (e.g., lacking lip-reading ability) or from the interference of audio input? To explore this, we conducted ablation experiments using the VideoLLaMA2 model series, chosen for its open-source availability of both vision-only and audio-visual variants.
As shown in the Table 5, VideoLLaMA2-7B-16F (vision-only) exhibits only a slight advantage over Video-LLaMA2.1-AV (audio-visual) on AVSM and ASD tasks, yet still lags far behind vision-only models such as VideoLLaMA3 and InternVL2.5 (Table 1). This indicates that VideoLLaMA2-7B has inherently poor lip-reading capability, which further implies that the audio-visual variant (Video-LLaMA2.1-AV) also suffers from limited visual lip-reading ability. Such limitations constrain its upper-bound performance in speech-visual alignment tasks. On the other hand, Video-LLaMA2.1-AV shows no significant performance advantage when utilizing audio information compared to its vision-only counterpart. This suggests that vocal information is not effectively leveraged, likely due to insufficient speech parsing capability in video MLLMs and inadequate understanding of cross-modal associations between audio and visual content.
9 Model Evaluation Implementation
Prompt. In order to facilitate the statistical model to answer the results, following common practices used in MLLM evaluations [24, 18], we adopt the following prompt to guide the MLLM to output option letters: “ Select the best answer to the following multiple-choice question based on the video. Respond with only the letter of the correct option. <Question-choices> Only answer best answer’s option letter. Your option is: ”. Evaluation Environments. All evaluation experiments for open-source models were conducted on a single NVIDIA L20 GPU with an inference batch size of 1.
Baseline Configurations and Runtime Statistics. Table 6 shows the scale, parameter settings, and costs (including memory usage and end-to-end testing time) for each model on HumanVBench. All hyperparameter settings follow the default configurations of these open-source works.
Model Time (min) top_p top_k num_beams temp. VRAM
Chat-UniVi (7B) 35 1 50 1 0.2 14G
CogVLM2-Video (8B) 34 0.1 1 1 0.2 26G
Video-LLaVA (7B) 46 1 50 1 1 37G
LLaVA-OneVision (7B) 46 1 50 1 1 33G
PLLaVA (7B) 35 0.9 50 1 1 18G
ShareGPT4Video (8B) 44 0.9 50 1 1 17G
Otter-V (7B) 60 1 50 3 1 17G
VideoChat2-IT (7B) 50 0.9 50 1 1 18G
InternVL2 (7B) 44 1 50 1 1.0 21G
InternVL2.5 (7B) 31 1 50 1 1.0 21G
Qwen2-VL (7B) 43 0.001 1 1 0.1 22G
Qwen2.5-VL (7B) 33 0.001 1 1 0.1 22G
LLaVA-Video (7B) 173 0.8 20 1 0.7 36G
Video-LLaMA3 (7B) 86 0.8 20 1 0.7 21G
Video-LLaMA (7B) 38 1 50 2 1 22G
Video-LLaMA2.1 (7B) 27 0.9 50 1 0.2 23G
ImageBind-LLM (7B) 75 1 50 1 1 22G
ChatBridge (13B) 27 1 50 1 0.2 28G
OneLLM (7B) 116 0.75 50 1 0.1 16G
GPT-4o 185 1 - - 1 API
Gemini-1.5-Pro 250 1 40 - 0.9 API
Table 6: Model configuration and runtime statistics evaluated on HumanVBench (one pass for all provided test samples).
10 Definitions and Examples for Each Task
10.1 Emotion Perception
Emotion Recognition aims to judge the overall emotional state of the person highlighted by a red bounding box in the video. An example is shown in Figure 7.
Figure 7: Example of Emotion Recognition task.
Emotion Temporal Analysis involves analyzing the changes in the emotions of the people highlighted with the red bounding box over time, identifying gradual intensification, diminishment, emotions shifts to test the model’s ability to track emotional dynamics. An example is shown in Figure 8.
Figure 8: Example of Emotion Temporal Analysis task.
Attitude Recognition involves inferring a character’s attitude towards things, categorized into four fixed options: positive, neutral, negative, and indeterminate. An example is shown in Figure 9.
Figure 9: Example of Attitude Recognition task.
Emotion Intensity Comparison requires compares the emotional intensity differences among various individuals in the video to find the most emotional person, assess whether the model can quantify and differentiate emotional intensity. An example is shown in Figure 10.
Figure 10: Example of Emotion Intensity Compare task.
10.2 Person Recognition
Text-to-Human requires the model to identify the specific person in a multi-person video based on a given text description, to test the model’s ability to locate and identify the described person. An example is shown in Figure 11.
Figure 11: Example of Text-to-Human task.
Human-to-Text asks the model to choose the most accurate description of the target person in a multi-person video, to ensure that the person is clearly distinguished from others and uniquely identified. This task requires the model to analyze and compare individuals in the video, identifying distinguishing features of the target person, such as appearance, clothing, actions, location, and other characteristics. An example is shown in Figure 12.
Figure 12: Example of Human-to-Text task.
Human Counting requires the model to determine the total number of distinct individuals in the video, testing its capability to detect, track, and accurately count individuals in complex scenes. An example is shown in Figure 13.
Figure 13: Example of Human Counting task.
Appearance Time Detection requires the model to identify the exact time frames when a specified person appears, demanding the ability to precisely mark the start time, end time, and duration of the individual’s presence in the video. An example is shown in Figure 14.
Figure 14: Example of Appearance Time Detection task.
10.3 Human Behavior Analysis
Behavior Temporal Analysis involves analyzing the dynamic changes in a specified person’s behavior over time, testing the model’s ability to accurately capture and track the temporal characteristics of these changes. An example is shown in Figure 15.
Figure 15: Example of Behavoir Temporal Analysis task.
Behavior Causality Analysis aims to investigate the causal relationships underlying a specific behavior, requiring the model to determine whether a person’s behavior in the video is triggered by a particular event or leads to subsequent actions. An example is shown in Figure 16.
Figure 16: Example of Behavior Causualty Analysis task.
Action at Specified Time asks the model to identify a person’s behavior or state at a specific time, testing its ability to accurately determine the person’s action or state at the given moment. An example is shown in Figure 17.
Figure 17: Example of Action at Specific Time task.
Time of Specific Action focuses on determining the time when a specific behavior occurs, requiring the model to accurately pinpoint the time of a particular action in the video. An example is shown in Figure 18.
Figure 18: Example of Time of Specific Action task.
10.4 Cross-Modal Speech-Visual Alignment
involves analyzing audio cues in multi-person videos to identify the individual whose appearance matches the voice. This task evaluates whether the model can recognize the voice gender and age and compare them with the appearance of the person in the video. An example is shown in Figure 19.
Figure 19: Example of Audio-Visual Speaker Matching task.
Active Speaker Detection asks the model to identify the active speaker in the video, requiring the model to accurately identify who is speaking by combining audio cues with the characters’ lip movements. An example is shown in Figure 20.
Figure 20: Example of Active Speaker Detection task.
Audio-Visual Alignment Detection requires detecting when the audio and video are synchronized, evaluating the model’s ability to synchronize audio and visual content, particularly through analyzing the speaker’s lip movements and voice. An example is shown in Figure 21.
Figure 21: Example of Audio-Visual Alignment Detection task.
Speech Content Matching requires matching the speech content of the video with text, validating the model’s ability to transcribe speech or translate lip movements into text. Figure 22 shows an example.
Figure 22: Example of Speech Content Matching task.
11 Annotations Details and Examples in Human-Centric Annotation Pipeline
For the in-the-wild videos collected from Pexels, we first apply splitting and filtering operations. Specifically, we begin by utilizing the video_resolution_filter, video_aesthetics_filter, and video_nsfw_filter operators to select videos that meet the following criteria: a resolution of at least 1280 in width and 480 in height, acceptable aesthetics, and appropriate content. Next, the video_split_by_scene_mapper is used to split the videos into scenes. The resulting clips are then filtered using the video_duration_filter to exclude clips shorter than 1 second and the video_motion_score_filter to remove static videos. These steps utilize existing operators in Data-Juicer [7], with parameters set to their default values except for the video_motion_score_filter, where the minimum motion score is set to 1.2. After completing these foundational steps, we apply the video_face_ratio_filter with a threshold of 0.65 to retain videos containing people. These videos are then processed using a series of mappers to generate fine-grained, multi-modal, human-related annotations.
We use a video example to demonstrate the annotation process and results, as shown in Figure 23. Below, we detail the models and settings used for each operator.
For the video_human_tracks_extraction_mapper, we follow the approach of Light_ASD [33], utilizing S3FD [59] as the face detector. A face bounding box is added to a human track if its overlap rate exceeds 50%. After obtaining the face track, we identify the corresponding body bounding box for each face bounding box in the same frame to generate a second bounding box track for the individual, referred to as the body track. The matching criterion selects the candidate bounding box with the smallest horizontal center distance and a smaller area. This process can be expressed by the following formula:
$\begin{split}&\text{closest\_bbox}=\mathop{\arg\min}_{\text{bbox}\in\text{candidate\_bboxes}}\\ &\left(\left(\frac{x_{1}+x_{2}}{2}-\frac{\text{f\_x1}+\text{f\_x2}}{2}\right)^{2}+(x_{2}-x_{1})(y_{2}-y_{1})\right),\end{split}$ (1)
where f_x1,f_x2 are the left and right boundaries of the face bounding box, the candidate human bounding boxes are obtained using YOLOv8-human, $x_{1}$, $x_{2}$, $y_{1}$, $y_{2}$ are the boundary values of a candidate human bounding box. If no bounding box meets the criteria, the frame is skipped, and detection proceeds with the other frames. Finally, the empty elements in the body track are replaced with the average of the bounding boxes from the surrounding frames.
In the human_demographics_mapper, we use DeepFace to perform frame-level detection of facial gender, age, and race. The analysis is conducted on cropped frames obtained directly from the video_human_tracks_extraction_mapper results. Finally, for a given face track, the demographics features are determined by taking the mode of the frame-level gender and ethnicity detections, and the median of the age detections.
In the video_human_description_mapper, we use the body bounding box track to crop the video, creating a reconstructed video focused on a single individual. This reconstructed video is then processed using ShareGPT4Video [11] for appearance description and simple actions.
In the video_facial_description_mapper, we use the face bounding box track to crop the video, creating face-focused reconstructed videos for emotion description using VideoLLaMA2.1 [16]. The choice of VideoLLaMA2.1 is based on a comparative analysis of multiple models, which revealed that VideoLLaMA2.1 is often more effective at identifying negative emotions. This capability is particularly important for adjusting emotion distribution before designing emotion recognition tasks.
The audio_tagging_mapper is a built-in operator in Data-Juicer, which we use directly for audio type classification.
The core model for the operator active_speaker_detection_mapper is ASD-Light [33]. Each face track sequence is analyzed together with the corresponding audio segment for the same time period. The model outputs a score sequence of the same length as the face track’s frames, where each score evaluates whether the individual is speaking in the current frame. Positive scores indicate active speaking, while negative scores indicate not. To assign a binary “speak or not” label to a human track, we classify an individual as an active speaker if the longest sequence of consecutive positive scores exceeds 12 frames. Notably, to reduce false positives, we cross-check the voice-based gender and age attributes with the individual’s demographic features. If there is a significant mismatch, the positive label is reassigned as negative.
The automatic speech recognition model used in the ASR_mapper is SenseVoice [1], which can also be utilized in the speech_emotion_recognition_mapper. For the voice_demographics_mapper, we use the wav2vec2 [2] model.
The results of the video_description_mapper are not directly involved in the construction of multiple-choice questions in this work. However, the environment, atmosphere, and events occurring in the video play a crucial role in understanding the actions and expressions of individuals. Therefore, we have included this mapper in the Human-Centric Annotation Pipeline. The example shown in Figure 23 is generated by ShareGPT4Video.
Notably, in the Human-Centric Video Annotation Pipeline, all the models we use are based on the most advanced open-source models available. As more powerful and specialized models emerge, integrating them into our pipeline can further enhance the quality of annotations.
12 Complete Construction Details of All Tasks
We will first explain the details of six descriptive questions generated using the Distractor-Included QA Generation Pipeline, followed by the construction details of the remaining tasks.
12.1 Construction Details of 6 Descriptive Human-Centric Questions
For these six tasks, the video-MLLM used to obtain task-oriented captions is VideoLLaMA-2.1[16]. The LLM used for generating question and initial answer pairs is Qwen2.5[54]. The three video-MLLMs employed for optimizing answers and producing raw distractors are VideoLLaMA-2.1[16], CogVLM2[21], and LLaVA-OneVision[28], respectively. The LLM responsible for generating distractors is Qwen2.5. We first present the general instruction templates in the six task generation processes.
The prompt template for the three Video-MLLMs used to refine answers and generate raw distractors is:
Please focus on the people whose heads are highlighted with red bounding boxes in the video and answer my question: ⟨Question⟩; Provide a brief response in one sentence.
The instruction used to compare the answers is:
Based on the video and my question: ⟨Question⟩Tell me which answer is better: (A). ⟨current model’s answer. ⟩(B). ⟨previous best answer⟩. Just answer (A) or (B).
The prompt template used in the “LLM for Generating Distractors” in Figure 3 is:
Below is a ready-made question and its multiple-choice options: ⟨Question⟩, Proper Answer: ⟨Answer⟩, Distractors1: ⟨eliminator1⟩, Distractors2: ⟨eliminator2⟩, Distractors3: ⟨eliminator3⟩. This question-option set may have the following issue: The current distractors have no errors; they simply represent alternative answers to the question. This makes the correct answer less distinct compared to the distractors. Therefore, I would like your help to add minor, distinct errors to each distractor so that the correct answer is clearly the only Proper Answer. Here are the minor errors type available for selection: ⟨error type⟩.
Remember that the modified distractors must meet the following requirements: 1. Be modified from the original distractor with only slight changes. You are not allow to creat new ones from scratch. 2. Be distinctly different from the Answer, without being overly semantically similar. Minor errors can be added. 3. Differ from each other. 4. Distractors should have similar length to the correct answer. If it is too short, lengthen the description.
In addition to the questions and options, the differences include the ⟨error types⟩. Next, we describe the construction of each task in detail.
Emotion Recognition: Since Label-5 is naturally a description based on face-focused cropping videos, it is directly used as the task-oriented caption. Additionally, Label-4 is included in the task-oriented caption to enhance the detail of the questions. Considering that most in-the-wild videos exhibit positive or neutral emotions, while we aim to ensure a sufficient proportion of negative emotions in the evaluation, videos for question generation are preselected at a ratio of positive:neutral:negative = 1:1:2. This selection is achieved by using an LLM to classify the emotional polarity of the descriptions. The resulting balanced category captions are used for question generation, with the following prompt:
Please generate one question and answer pair based on the person’s description: ⟨task-oriented caption⟩. the question should closely related to emotion recognition. Here is an question example: “What emotions might the girl in red dress be experiencing during her practice?”
The video for the question is marked using the face bounding box from the target character’s Label-1. The type of minor errors (⟨error types⟩) introduced for building distractors is: Add incorrect emotional descriptors or modify the original emotional descriptors to incorrect ones.
Emotion Temporal Analysis: We first select videos longer than 7 seconds and then identify described characters with emotional changes based on Label-5 (using an LLM for binary classification). The videos of these characters are used for question generation. For this task, Label-5 is directly used as the task-oriented caption, with Label-4 added to enhance the details of the questions. The question generation prompt is:
Please generate a question-answer pair based on the following video caption. Please note that the questions must be related to the emotional temporal changes. Here are some example: 1. How does the girl in red’s emotions change as the video progresses? 2. How does the girl in red’s emotions change as she dances in the video?
The video for the question is marked using the face bounding box from the target character’s Label-1. The type of minor errors introduced for building distractors is: Add some incorrect emotions to the sequence, remove some correct emotion words, or change the original emotional descriptors to incorrect ones.
Behavior Temporal Analysis: First, videos longer than 7 seconds are selected for question generation. Then, the target character in the video is highlighted using the face bounding box track from Label-1. Based on the marked videos, appearance cues of the target character (i.e., Label-4) are added to help to guide the model’s attention to the individual. The prompt for obtaining the task-oriented caption is designed as follows:
Please focus on the person highlighted by the red bounding box (⟨Human_Appearance⟩) and tell me if the actions of the person changed over time and what actions does the person take in order? Respond according to the following format: {“Action_Change”: True or False, “Action_Sequence”: action sequence}.
Based on the task-oriented captions, select characters with changes in actions for LLM question generation. The prompt for question generation is: Please generate a question-answer pair based on the following human’s behavior caption. The generated questions should focus on identifying the action sequence of the highlighted person. The following is a description of a human in the video. ⟨action_sequence⟩. The focused person is ⟨appearance⟩. Here is a question template you can refer to: What actions and behaviors does the girl in the red dress display in the video in order? List them sequentially. Remember do not reveal the answers in your questions and the answer should be brief and just in one sentence.
The type of minor errors introduced for building distractors is: Add some nonexistent actions, remove some actions, or replace correct actions with incorrect ones.
Emotion Intensity Compare: First, count the number of frames corresponding to the track with the longest appearance time in each video. If there are more than three and less than seven tracks in the video that reach this number of frames, keep the video for question generation. This step mainly use the information from Label-1 . All individuals corresponding to the tracks with the most frames will be used for question generation. The question video is created by utilizing these human tracks to mark the individuals and adding letter labels. For this task, initial question-answer pairs are directly created. The question is, “Which person in the image, ⟨LETTERS⟩, has the strongest emotion? What specific emotion is it? Please respond briefly in the format ⟨letter-emotion⟩.”, in which ⟨LETTERS⟩ refers to all the selectable individuals’ letter labels. The answer is, “The emotional intensity of the selectable characters in the image is similar, and they are all neutral.” The subsequent three models will refine the answer.
The type of minor errors introduced for building distractors is: If the letters are the same, minor modifications to the emotions can be made to make the options different; if the letters referring to people are different, the emotions can remain unchanged.
Human-to-Text: First, select videos with 3 to 7 individuals based on Label-2, and then choose the person who appears the most frames in the video as the target individual for question generation. Next, highlight the target individual in the video using the face bounding box track from Label-1. Based on the marked video, appearance cues of the target individual (i.e., Label-4) are added to help the model focus on the person. The prompt for obtaining the task-oriented caption is designed as follows:
Please accurately describe the person highlighted by a red box(⟨appearance⟩), your answer can be based on appearance, location, and actions, so that the highlighted person can be distinguished from others in the video. Please respond in only one sentence and begin with “The person is …”.
Based on the above description of the target individual, the question-answer pair is directly constructed. The question is fixed as: “Please accurately describe the person highlighted by a red box based on appearance, location, and actions, so that the highlighted person can be distinguished from others in the video.” The initial answer is the task-oriented caption.
The type of minor errors introduced for building distractors is:Based on the items, people, and position information, add small modifications to make the location information incorrect; alternatively, you can also modify the description of the person’s appearance to introduce errors.
Behavioral Causality Analysis: The construction process is similar to the design process of Behavior Temporal Analysis. First, videos longer than 7 seconds are selected for question generation. Then, the target individual in the video is highlighted using the face bounding box track from Label-1. Based on the annotated video, appearance cues of the target individual (i.e., Label-4) are added to assist the model in identifying the person to focus on. The prompt for obtaining the task-oriented caption is designed as follows:
Please describe the causal events related to the person highlighted by the red bounding box (⟨appearance⟩) in the video: what causes this person to exhibit a certain behavior, or what actions does this person take that led to a certain event. If no causal events exist, respond without causal events. Please answer in the following format: {“causal_events_exist”: True or False, “causal_events_description”: description}.
Videos and target individuals with causal relationships (i.e. “causal_events_exist” is true) are then selected for question generation. The prompt for question generation is: Please generate a question-answer pair based on the following video caption. The generated questions should inquire about causal reasoning related to the character’s expressions or behaviors. The following is a description of the human in the video: ⟨causal_events_description⟩. The focused person is ⟨appearance⟩. You should either follow the causal analysis question template “Analyze why the girl in the red dress raises her hand.” or the result derivation question template “What does the girl in the red dress raising her hand lead to?”. Remember do not reveal the answers in your questions and the answer should be brief and just in one sentence.
The type of minor errors introduced for building distractors is: Explain the result using incorrect causes, misdescribe the effect of the cause-and-effect relationship, reverse the order of cause and effect, exaggerate or minimize factors.
12.2 Construction Details of 10 Closed-Ended Human-Centric Questions
Attitude Recognition: This task is constructed based on the first half of the Distractor-Included QA Synthesis Pipeline. The human who appears in the most frames is selected as the target for question generation. The target individual is highlighted in the video using the face bounding box track from Label-1. Based on the annotated video, appearance cues of the target individual (i.e., Label-4) are added to the prompt to help the model focus on the intended person. The prompt used to obtain the task-specific caption is:
Focus on the person highlighted by the red bounding box (⟨appearance⟩) and tell me: Do the highlighted people display certain attitudes toward specific objects and events? What kind of attitude is it?
The prompt used for question generation is:
Please generate a best question-answer pair based on the following video caption. The generated questions should focus on analyzing the character’s attitude, which should be one of positive, negative, or neutral. The following is a description of the human in the video. ⟨task-specific caption⟩The focused person is ⟨appearance⟩. Here are some question templates you can refer to: 1. What is the attitude of the girl in the blue shirt towards taking the bus in the video? positive, negative, or neutral? 2. What is the woman in the beige jacket’s attitude? Positive, negative, neutral? Please remember not to reveal the answers in your questions and the answer should be brief and just in one sentence.
The options consist of four choices: Positive, Negative, Neutral, and Indeterminate. The Indeterminate option is included as a supplemental choice to ensure answers optional.
Text-to-Human: The criteria for selecting the videos for questioning are consistent with the Emotion Intensity Compare task selection rules. Then, use the same method as Human-to-Text to obtain task-specific captions and directly use the description of the target person to complete the question template: “Please select the person in the video that best matches the following description: ⟨human description ⟩”. The video corresponding to the question is marked with the face bounding boxes of all individuals using Label-1, and each individual is distinguished by a capital letter label. The selectable options are the letter labels representing each person.
Human Counting: For an annotated video, the approximate number of people in the video can be estimated directly using Label-2. However, due to issues such as blurred crowd background, overlapping between people and objects and other factors, this estimate is often imprecise, especially in crowded scenes. Therefore, Label-2 is only used to adjust the question distribution (3–5 people: 60, 6–8 people: 60, 9+: 54). The ground truth number is manually annotated, and distractors are constructed based on this value. The distractors construction rule is to randomly select three different numbers within a range of up to 4 from the ground truth number, excluding the ground truth itself.
Appearance Time Detection: First, select videos based on the following criteria: the video duration must exceed 7 seconds, and the target individual’s presence should account for between one-third and two-thirds of the total video length (calculated as the ratio of the human track frames to the total frames), primarily using Label-1. Then the frame range from Label-1 is used to determine the target individual’s appearance time range (format both ends as integers), which serves as the ground truth for generating questions about this person.
To obtain a detailed and accurate description of the individual, the same method as in the Human-to-Text task is used to generate the task-specific caption for the target. Using the description, questions are constructed in a template-based manner, as shown in Figure 14.
For distractor construction, three random time intervals are generated near the ground truth time interval, ensuring that their overlap with the ground truth interval does not exceed 4 seconds. This ensures the distractors do not cause confusion when selecting the correct answer.
Note that in this task, videos with bounding boxes are only used during the automatic description generation by Video-MLLMs and for manual verification. In the final version of the questions, the videos do not include bounding boxes.
Action at Specified Time and Time of Specific Action tasks rely on manual annotation, as attempts with various open-source models revealed their inability to accurately identify the timing of specified actions. For manual annotation, annotators are required to watch the videos and observe whether the highlighted individual performs any distinct short-term actions (quickly completed actions or “the start of an action”, but not continuous states). They should record the action and its starting time. The videos and target individuals are consistent with those in Behavior Temporal Analysis task. Based on the specific action–time pairs provided by the annotators, two types of action-time-related questions are constructed.
Human Emotion Perception Person Recognition Human Behavior Analysis Speech-Visual Alignment
Labels ER ETA AR EIC T2H H2T HC ATD BTA BCA AST TSA AVSM ASD AVAD SCM
Label-1
Label-2
Label-3
Label-4
Label-5
Label-6
Label-7
Label-8
Label-10
Table 7: Annotation labels used in the construction process of 16 tasks
For the Action at Specified Time task, the question video consists of the highlighted target individuals with red bounding boxes. Only video samples where short-term actions are present are selected for question generation. The question template is shown in Figure 17. The ground truth is the specific action annotated for the individual. Distractors are generated by LLM from the task-specific captions in Behavior Temporal Analysis, with the following prompt:
Please select and modify 3 actions from the list below to ensure that each action is significantly different from the target action ⟨ground truth action⟩. Here is the original action list: ⟨action_list⟩. Begin with “begin to ..” for an action. If the number of actions is less than 3, generate one.,
where the ⟨action_list⟩is the sequence of actions generated by using LLM to summarize the task-oriented caption.
For the Time of Specific Action task, the question video is marked with target individual’s bounding boxes. The question template is shown in Figure 18. For question samples with short-term actions, distractors are generated by selecting three numbers that are at least 3 seconds apart from the ground truth action time. For videos where a continuous action state is maintained throughout, the ground truth is set to “This is an ongoing action in the video.” The distractors for such cases are fixed at 1s, 4s, 7s, and 10s. Note that to keep the options consistent, each question includes the option “This is an ongoing action in the video.”
Audio-Visual Speaker Matching: First, select the appropriate question videos. The constraints mainly include video conditions and character conditions. The video conditions include: the audio label being “Speech”, the number of people in the video is between 2 and 4, and the video duration is not less than 4 seconds; the character condition is: the frame coverage of the character must reach more than 67% of the video frame number. Further, the target person is selected as the ground truth according to the correlation between the age and gender attributes of the audio and the appearance of the person. Specifically, if the audio age belongs to “child”, the only child is selected from the video as the target person, and the other characters are interference characters; If the audio age is “adult”, the only adult with the same gender as the audio is selected from the video as the target person, and the other characters are interference characters. The age information is binary here because the audio age attribute is relatively vague. In addition, the gender characteristics of children’s voices are not always distinguishable. Therefore, in order to enhance the optionality of the answer, the character types are divided into only three categories: male, female, and child. The question video uses Label-1 to mark each optional person and capital letters as the option.
Active Speaker Detection: Suitable question videos are selected based on the criteria of having an audio label of “Speech”, 2 to 4 people, and a duration of at least 4 seconds. The video must contain a single active speaker. Label-1 is used to label all individuals, with the active speaker’s label as the ground truth and others as distractors. Since the automated active speaker labels may not always be reliable, two additional options are included for each question to facilitate manual correction later: “There are more than one active speaker in the video.” and “There is no active speaker in the video.”
Audio-Visual Alignment Detection: Suitable question videos are selected based on the criteria of having fewer than 3 people, a duration of over 8 seconds, an audio type of “speech”, and at least one active speaker. The video is then divided into three equal segments, with the left endpoint of each segment (rounded to an integer) used as potential options. One of these options is randomly chosen as the ground truth. The video is then modified by reversing the audio before the selected timestamp to create a “misaligned audio-visual” video.
Speech Content Matching: Videos are selected for question creation based on the following criteria: single-person scenes, duration greater than 5s, audio type “speech”, and the person is an active speaker, the speech content being English with its sentence length greater than 35 characters. The ground truth is the automatic speech recognition result corresponding to Label-8. Distractors are generated using an LLM, which creates three different sentences with similar meaning and length to the ground truth as distractors.
In Table 7, we illustrate which labels from the Human-Centric Video Annotation Pipeline are used to construct each task.
12.3 Details of Human Efforts for HumanVBench
We employed two professional full-time AI data annotators and invited two graduated-level volunteers to participate in the construction and evaluation of HumanVBench. They collaborate to complete a series of tasks, with two main annotators participating in the full data of each task and two volunteers participating in the sampled data. Inconsistencies will be identified and resolved (for example, through discussion or majority voting to reach a consensus) to ensure high quality. The average annotation time per question is 3 minutes, totaling 10 workdays. Specifically, their tasks include the following four parts.
Human Annotation on Generated QAs: Tasks requiring human annotations to generate questions and options include Human Counting, Action at Specified Time, and Time of Specific Action. The latter two tasks can streamline annotation by annotating a single dataset containing “special short-term action & moment of occurrence”. Therefore, in this step, each annotator was assigned to one annotation set. All other tasks were generated automatically, reducing the cost of human annotation.
Manual Verification and Correction: Except for the tasks that are already reliable enough, which include: Emotion Recognition in Conversation task reconstructed from the MELD benchmark, the three tasks derived from the aforementioned human annotations, and the Audio-Visual Alignment Detection and Speech Content Matching tasks, all other tasks require manual verification to ensure quality, following the process described in Section 3.3.4. During correction, low-quality samples (e.g., person transitions in human tracking, video freezing midway) are required to be flagged for removal.
Cross-Verification: After completing the above steps, we obtained 16 usable tasks for evaluation. To further ensure the high quality of the questions, we conducted cross-verification on 16 tasks except Emotion Recognition in Conversaton task to reduce the impact of personal biases and errors on the benchmark. Specifically, the tasks were cross-assigned to the two major annotators. For each task, the annotator in this step was ensured to be different from those responsible for the Manual Verification and Correction or Human Annotation. Annotators were first required to answer the multiple-choice questions. For disputed questions where answers were marked “incorrect”, the new correct answer or option will be updated for this question, and these disputed questions were reassigned to another annotator for a second review. If errors persisted, both annotators discussed and agreed on a unified answer to serve as the final ground truth.
Human-Evaluation Baseline in Table 1: For each task except the Emotion Recognition in Conversation task, we define the number of correct human responses based on the cross-checking process as follows: the sum of the number of questions answered correctly by the first annotator after completing all questions and the number of questions answered correctly by the second annotator on the disputed questions after correction by the first annotator. The human response accuracy for the task is calculated as the ratio of the number of correct human responses to the total number of questions in the task.
Figure 23: An example of using Human-Centric Annotation Pipeline for annotation.